OpenStudy (anonymous):

x^2 + 2xy + (x^2 - 1/3x^3 - x^2y)dy/dx = 0 Solve this DE by finding a function of the form u = u(y), as an integrating factor. Seperation of variables ... then ?

OpenStudy (turingtest):

no separation of variables, you need to get this into the form\[y'+p(x)y=q(x)\]then the integrating factor will be\[\mu=\exp\left(\int p(x)dx\right)\]

OpenStudy (turingtest):

wait, the IF is supposed to be a function of y ???

OpenStudy (anonymous):

ya ...

OpenStudy (anonymous):

so i'm kind of confused.